OpenStudy (hang254):

A=1/2h (b1 + b2) for b2

OpenStudy (anonymous):

whats your Q????

OpenStudy (hang254):

the one above

OpenStudy (anonymous):

you need numbers or something

OpenStudy (anonymous):

?????

OpenStudy (hang254):

all it says is solve for the indicated variable. which appears to be b2

OpenStudy (anonymous):

\[A=\frac{ 1 }{ 2h } (b1 + b2) \]

OpenStudy (anonymous):

ok just to be clear i will multiply \[\frac{ 1 }{ 2h }\] by (b1+b2) we have A=b1/2h + b2/2h

OpenStudy (hang254):

ok, i am following you

OpenStudy (anonymous):

to get rid of 2h on hte right side we multiply by 2h so >>> A(2h)=(2h)(b1/2h + b2/2h)

OpenStudy (anonymous):

simplifying the equation we get>>>>> 2hA=b1+b2

OpenStudy (anonymous):

are we clear??

OpenStudy (hang254):

yes, i am reading what u wrote right now

OpenStudy (anonymous):

good

OpenStudy (hang254):

thanks

OpenStudy (anonymous):

and finally 2hA-b1=b2.

OpenStudy (hang254):

thank you

OpenStudy (anonymous):

your most welcome :)

OpenStudy (anonymous):

try to think what operators you can use to get rid of fractions so you can simplify it.

OpenStudy (hang254):

ok. i will, thanks again